Which two functions are inverses of each other. Show that they are inverse functions. f(x)= x+2 over 4, g(x)= 4x+2, h(x)= x-2 over 4 .
Answer by stanbon(48558)

g(x)= 4x+2, h(x)= x-2 over 4
-----------
g and h are inverse to one another:
g[h(x)] = g[(x-2)/4] = 4[(x-2)/4] + 2 = x-2+2 = x
h[g(x)] = h[4x+2] = [4x+2 - 2/4 = 4x/4 = x
===============
Cheers,
San H.